Given f(x) = 3x3 + 4x2 + 4x + 3. Find (-1)0).
Answer #1

1 1 Solution Given /(x) = 3x +4x2 +4x+3 (11)(0) We have, Formula : (s)(x)= 7*(7) We need to find : (1-1)0) = (s)(0-7410) C